Which is the best estimate for 3 1/3+6 3/4

Answer by oberobic(2304)

3 1/3 = 10/3 = 40/12
.
6 3/4 = 27/4 = 81/12
.
40/12 + 81/12 = 121/12
.
121/12 = 10 1/12